The car engine does the job 4200kJ in 1 minute. Calculate its power in kW.


1
Expert's answer
2021-03-31T09:40:16-0400

By the definition of the power, we have:


"P=\dfrac{W}{t}=\dfrac{4200\cdot10^3\ J}{60\ s}=70\ kW."